OpenStudy (gabby_taylor1224):

MEDAL! a circle has its center at (2 -4) and a radius of 6 units. what is the equation of the circle

OpenStudy (bemoseemo):

what kind of equation? if you need an equation for the circumference the answer would be \[\pi*r ^{2}\]

OpenStudy (bemoseemo):

but based on your information, do you need the equation for the location of the verticies or...

OpenStudy (gabby_taylor1224):

Im not sure thats all the question says

OpenStudy (gabby_taylor1224):

I would probably assume the circumference

OpenStudy (bemoseemo):

here ya go x2 + y2 - 4x + 8y - 16 = 0 yeesh that took a while

OpenStudy (gabby_taylor1224):

so would it be 3.14 X R ^2 what would be my radius?

OpenStudy (gabby_taylor1224):

so x2 + y2 - 4x + 8y - 16 = 0 would be my answer?

OpenStudy (bemoseemo):

yes

OpenStudy (bemoseemo):

my brain needs to rest after a long day of questions like these

OpenStudy (gabby_taylor1224):

thank you so much!

OpenStudy (bemoseemo):

yw tag me if you need any more help but I am going to go for a bit